Q4. If you wish to produce 45.0 nm x-rays in the laboratory, what is the minimum voltage you must use in accelerating the electrons?
1
Expert's answer
2020-06-08T10:30:03-0400

In our case the minimum voltage is the photon energy with the wavelength of 45.0 nm:


"E_{photon} = \dfrac{hc}{\lambda},""E_{photon} = \dfrac{6.626 \cdot 10^{-34} \ J \cdot s \cdot 3 \cdot 10^8 \ \dfrac{m}{s}}{45 \ nm},""E_{photon} = \dfrac{1.9878 \cdot 10^{-25} \ J \cdot m \cdot \dfrac{1 \ eV}{1.602 \cdot 10^{-19} \ J}}{45 \ nm},""E_{photon} = \dfrac{1240 \cdot 10^{-9} \ eV \cdot m \cdot \dfrac{10^9 \ nm}{m}}{45 \ nm},""E_{photon} = \dfrac{1240 \ eV \cdot nm}{45 \ nm} = 27.5 \ eV."

Answer:

"E_{photon} = 27.5 \ eV."